How would I solve this? The gravitational force between two objects varies inversely as the square of the distance between the objects. If a force of 25 pounds results from two objects that are 6 miles apart, how much force results from two objects that are 15 miles apart?
Answer by Alan3354(69443)

The gravitational force between two objects varies inversely as the square of the distance between the objects. If a force of 25 pounds results from two objects that are 6 miles apart, how much force results from two objects that are 15 miles apart?
-----------------
F = k/r^2 (r is the distance)
25 = k/6^2
k = 900
-----------
F = 900/15^2 = 900/225
F = 4 pounds
-----------------------
It's not necessary to solve for k:
15 = 2.5*6
F = 25/(2.5^2) = 25/6.25 (inverse square)
F = 4 pounds